OpenStudy (anonymous):

I need help with understanding a proving trigonometric equations problem: -tan^2 x+ sec^2 x =1

OpenStudy (anonymous):

couldn't tell you

OpenStudy (anonymous):

Would it be the same as sec^2 x - tan^2 x =1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're in 7th I learned this last year but i dont remember this srry

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos ^2 }-\frac{ \sin^2 }{ \cos ^2 }\]

OpenStudy (anonymous):

i figured it out

OpenStudy (anonymous):

\[\frac{ 1-\sin ^2 }{ \cos ^2 }=1\]

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

\[1-\sin ^2=\cos ^2\]

OpenStudy (anonymous):

\[1=\cos ^2+\]

OpenStudy (anonymous):

sin^2

OpenStudy (anonymous):

ok you can follow that my CPU is lagging

OpenStudy (anonymous):

or can you

OpenStudy (anonymous):

I'm sorry, I'm still confused. Shouldn't 1-sin^2/cos^2=1 be the final solution?